OpenStudy (anonymous):

plz help me!!! A survey of 594 students about the sports program at a school finds the following results: 495 students like the program. 99 students think the program is inadequate. If a circle graph were made from the data, what would the measure of the central angle be for the group that thinks the program is inadequate?

OpenStudy (anonymous):

I'm working on it. Hold on.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Is there any other information I have to go on? I don't want to fudge the math.

OpenStudy (anonymous):

I'm about to solve but, I don't think I have enough information.

OpenStudy (anonymous):

oh i forgot to put the answer's in here they are A.17 B.60 C.83 D.300 that is it

OpenStudy (anonymous):

Got it! Now let me explain.

OpenStudy (anonymous):

There are 594 students total. You said 495 agree and 99 disagree. So we need to find the percentages. 495/594 = 83.3% and 99/594 = 16.7% Now we can move on to converting it to angles. (495/594) x 360 = 178200/594 = 300degrees (99/594) x 360 = 35640/594 = 60 degrees This makes sense because there is 360 degrees in a circle. But we want the angle of the people who don't want the program, right? That would be 60 degrees. So the answer is B, 60 degrees. Hope that helps! :)

OpenStudy (anonymous):

Thank's ;p

OpenStudy (anonymous):

BTW, can I ask you a question? I wouldn't happen to know you in real life would I? Your username reminds me of someone...

OpenStudy (anonymous):

my name is faith smith i'm 13

OpenStudy (anonymous):

you can look me up on facebook

OpenStudy (anonymous):

Oh, then I don't think I know you. But I might friend you.

OpenStudy (anonymous):

okay ;)